zarasht:

A reaction has a rate law of Rate = (1.25 M⁻²s⁻¹)[A]². What concentration of would give the reaction a rate of 0.0495 if the concentration of [A] is 0.250 M?

justjm:

It seems like the rate law is, if I'm not wrong, \(r=k\left[A\right]\left[B\right]^{2}\) where \(k=1.25\ M^{-2}s^{-1}\) You are given that at a specific time, \(\left[A\right]=0.250\ M\) yielding a rxn rate of \(r=0.0495\) To find \(\left[B\right]^{2}\), you can just substitute the values and solve. \(0.0495=\left(1.25\right)\left(0.250\right)\left[B\right]^{2}\)

zarasht:

But it is asking to find just B

Mercury:

once you have B^2 you can simply take the square root to find B